OpenStudy (anonymous):

integrate from 0 to (pi/2) ((1/2)sin1)d theta please help integrate from 0 to (pi/2) ((1/2)sin1)d theta please help @Mathematics

OpenStudy (anonymous):

is this really \[\int_0^{\frac{\pi}{2}}\frac{1}{2}\sin(1)d\theta\]???

OpenStudy (anonymous):

I got pi/4 sin 1 from Wolfram alpha but I don't know how it got that?

OpenStudy (anonymous):

is it really sine of 1?? that is a constant. seems odd

OpenStudy (anonymous):

yes sin1

OpenStudy (anonymous):

\[\sin(1)\] is just a number. you are integrating a constant from 0 to pi/2

OpenStudy (anonymous):



OpenStudy (anonymous):

just have a rectangle with base \[\frac{\pi}{2}\] and height \[\frac{1}{2}\sin(1)\] so area is \[\frac{\pi}{2}\times \frac{1}{2}\sin(1)\]

OpenStudy (anonymous):

so I just "add" pi and plug pi/2 in?

OpenStudy (anonymous):

there is no "adding" it is base times height. base is pi /2, height is 1/2 sin(1)

OpenStudy (anonymous):

if you look at your question it says integrate with respect to theta, but there is no theta there. you just have a constant

OpenStudy (anonymous):

why is it different than integrating 1 dx which turns into x then you evaluate x at a and b? I haven't done integration for a long time.

OpenStudy (anonymous):

like \[\int_0^5 4 dx=5\times 4\]

OpenStudy (anonymous):

you could do the same thing and you will get the same answer.

OpenStudy (anonymous):

oh, so, using your example, I'd have 4x evaluated at 5 and 0 right?

OpenStudy (anonymous):

but it is silly to integrate a constant using anti-derivatives. you are trying to find the area of a rectangle, use base times height

OpenStudy (anonymous):

I'll use the rectangle from now on.

OpenStudy (anonymous):

yes in my example you could say the antiderivative of 4 is 4x, plug in 5, get 20, plug in 0, get 0, so answer is 20. but that is a waste of time

OpenStudy (anonymous):

a constant is a horizontal line, so you are finding the area of a rectangle. base time height does it. in other words it is the constant times the length of the path

OpenStudy (anonymous):

yes! thank you for pointing this out.

OpenStudy (anonymous):

yw